OpenStudy (anonymous):

PLEASE HELP FOR MEDAL + FAN!!! What is the slope of a line parallel to the line with equation 2x – 5y = 9?

OpenStudy (anonymous):

OPTION 1 \[-\frac{ 5 }{ 2 }\] OPTION 2 \[-\frac{ 9 }{ 2 }\] OPTION 3 \[-\frac{ 2 }{ 5 }\] OPTION 4 \[\frac{ 2 }{ 5 }\]

OpenStudy (blake57roger):

b

OpenStudy (blake57roger):

medal

OpenStudy (anonymous):

show proof its B

OpenStudy (anonymous):

well i found out is right so nvm ill giv u metal

OpenStudy (sleepyjess):

@blake57roger , Please do not give out direct answers.

OpenStudy (anonymous):

2x-5y=9 Put into y-intercept for and u get y=2/5x-9/5 The slope of a parallel line would b the same so it would be 2/5 @AMYCARTER @blake57roger and @Queen_Bee1 u might want to double check